Circular No. 7060

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998en IN UGC 3645 A. V. Filippenko and T. Matheson, University of California at Berkeley,

report that a CCD spectrum (340–1000 nm) obtained on Nov. 25 with the Shane 3-m reflector at Lick Observatory shows that SN 1998en (cf. IAUC 7045) is a type-II supernova. Broad Hα emission is visible, with less prominent P-Cyg absorption. The redshift of the object, determined from narrow emission lines from a superposed H II region, is 0.021.

RX PUPPIS R. J. Ivison, University College, London, writes: “Spectroscopy from

the 4.2-m William Herschel Telescope on Nov. 24 UT shows that the slow optical brightening of the symbiotic Mira-type variable star RX Pup has been accompanied by its return to a state of high excitation, with a forest of optical emission lines now visible. The Balmer Hα emission may be extended — tentative evidence of a mass-ejection event. Radio and x-ray monitoring is urged, as is multicolor photometry and long-slit spectroscopy.”

COMET P/1998 W1 (SPAHR)

Improved orbital elements from MPEC 1998-W35:

T = 1999 Jan. 18.858 TT  ω = 347.^{o}281
e = 0.51510  Ω = 101.819 }2000.0
q = 1.72642 AU  i = 22.077

a = 3.56033 AU n^{o} = 0.146713 P = 6.72 years

1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 23 4^{h}28.^{m}29 − 9^{o}36.^{′}6 0.887 1.806 148.^{o}9 16.^{o}4 15.8
Dec. 3 4 19.44 − 7 16.7 0.856 1.781 150.4 15.9 15.7
13 4 10.73 − 3 59.4 0.846 1.761 148.1 17.2 15.6
23 4 03.93 + 0 03.5 0.858 1.745 142.6 20.0 15.6
Jan. 2 4 00.44 + 4 33.1 0.890 1.734 135.5 23.4 15.6
12 4 01.01 + 9 09.6 0.940 1.728 127.8 26.7 15.7
22 4 05.82 +13 37.4 1.006 1.727 120.3 29.5 15.9

Total CCD magnitude and coma-diameter estimates: Nov. 20.25 UT, 15.8, 0′.34 (P. E. Roques, Williams, AZ, 0.25-m reflector + Kron-Cousins V filter); 23.16, 14.3, 0′.5 (T. Johnston, Boulder, CO, 0.45-m reflector).

1998 November 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green